A scientist working late at night in her low-temperature physics laboratory decides to have a cup of hot tea, but discovers the lab hot plate is broken. Not to be deterred, she puts about 8.00 oz of water, at 12.0°C, from the tap into a lab dewar (essentially a large thermos bottle) and begins shaking it up and down. With each shake the water is thrown up and falls back down a distance of 25.6 cm.

Is this practically possible?



1
Expert's answer
2022-05-28T08:23:09-0400

"Q_0=2mgh,"

"Q=nQ_0=cm\u2206t,"

"n=\frac{cm∆t}{2mgh}=\frac{c∆t}{2gh},"

"T=\frac n{T_0}=\frac{c∆t}{2ghT_0}=31.4^h."

Impossible.